OpenStudy (sharpsage):

My teacher didn't really explain this well and I just need to understand how to solve this. Tim Worker is doing his budget. He discovers that the average miscellaneous expense is $45.00 with a standard deviation of $16.00. What percent of his expenses in this category would he expect to fall between $38.60 and $57.80? The z for $38.60 = - ________ The percent of area associated with $38.60 =_____% The z for $57.80 =________ The percent of area associated with $57.80 =_____% Adding the two together, the percent of his expenses between $38.60 and $57.80 is _______%

Directrix (directrix):

Formula for z.

Directrix (directrix):

@sharpsage

Directrix (directrix):

I'll help you get started. The z for $38.60 = ? x = 38.60 z = 38.60 - 45 ____________ 16 z = - 6.40 ______ 16 z = - .40

Directrix (directrix):

The percent of area associated with $38.60 =_____% What about this? Do you have a table of areas under the normal curve in your book?

Directrix (directrix):

If you would help me help you, then we could knock out this problem. I need to know this to proceed but you won't tell me. >>Do you have a table of areas under the normal curve in your book?

OpenStudy (sharpsage):

Ok, Sorry I was writing everything down. I think I figured it out The z for $38.60 = is -0.4 and I for the percentage I got 15.54% and using the equation for z I got 0.8 for $57.80 and the percent of area associated with $57.80 is 28.81% I think So 44.35% would be the percent of his expenses between $38.60 and $57.80

OpenStudy (sharpsage):

@Directrix

Directrix (directrix):

What I am asking you is what you are using to get from a z of -.4 to a percentage. How do you know that >> for the percentage I got 15.54% I am not saying that it is wrong. I am asking how you got it.

OpenStudy (sharpsage):

Oh, sorry. The assignment included a table of areas. I didn't see it before.

Directrix (directrix):

Okay. That is what I was asking.

Directrix (directrix):

You would expect 44.356 per cent to fall between $38.60 and $57.80. Yes.